OpenStudy (sleepyjess):

Please help Divide f(x) by d(x), and write a summary statement in the form indicated. \(\dfrac{x^4-4x^3+2x^2-4x+1}{x^2+1}\)

OpenStudy (phi):

see http://www.khanacademy.org/math/algebra2/polynomial_and_rational/dividing_polynomials/v/polynomial-division also, write you bottom as x^2 + 0x + 1 when you do this problem

OpenStudy (sleepyjess):

Thank you!

OpenStudy (sleepyjess):

I came up with \(x^2-4x+1\) but it wants one of these as the answer \(f(x) = (x^2 + 1)( x^2 + 4x + 1)\) \(f(x) = (x^2 + 1)( x^2 - 4x + 1) + 12x + 3\) \(f(x) = (x^2 + 1)( x^2 + 4x + 1) + 12x + 3\) \(f(x) = (x^2 + 1)( x^2 - 4x + 1)\)

OpenStudy (sleepyjess):

@phi

OpenStudy (phi):

did it divide evenly?

OpenStudy (sleepyjess):

yes there was no remainder

OpenStudy (phi):

you found f(x) / d(x) = x^2−4x+1 if we multiply both sides by d(x) we get f(x) = d(x)(x^2−4x+1) d(x) is short for the denominator (x^2+1)

OpenStudy (phi):

refresh your browser if not clear.

OpenStudy (sleepyjess):

So it would be the last option?

OpenStudy (phi):

yes